OpenStudy (anonymous):

Evaluate 3 + 2 • 4. A. 20 B. 24 C. 11

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Multiplication takes precedence over addition in the order of operations, so 3 + 2*4 = 3 + 8 = 11.

OpenStudy (anonymous):

c. 2*4+3=11